### v2

# Benchmarking Spatiotemporal Reasoning in LLMs and Reasoning Models: Capabilities and Challenges

Abstract
Spatiotemporal reasoning plays a key role in Cyber-Physical Systems (CPS). Despite advances in Large Language Models (LLMs) and Large Reasoning Models (LRMs), their capacity to reason about complex spatiotemporal signals remains underexplored. This paper proposes a hierarchical SpatioTemporal reAsoning benchmaRK, STARK, to systematically evaluate LLMs across three levels of reasoning complexity: state estimation (e.g., predicting field variables, localizing and tracking events in space and time), spatiotemporal reasoning over states (e.g., inferring spatial and temporal relationships), and world-knowledge-aware reasoning that integrates contextual and domain knowledge (e.g., intent prediction, landmark-aware navigation). We curate 26 distinct spatiotemporal tasks with diverse sensor modalities, comprising 14,552 challenges where models answer directly or by Python Code Interpreter. Evaluating 3 LRMs and 8 LLMs, we find LLMs achieve limited success in tasks requiring geometric reasoning (e.g., multilateration or triangulation), particularly as complexity increases. Surprisingly, LRMs show robust performance across tasks with various levels of difficulty, often competing or surpassing traditional first-principle-based methods. Our results show that in reasoning tasks requiring world knowledge, the performance gap between LLMs and LRMs narrows, with some LLMs even surpassing LRMs. However, the LRM o3 model continues to achieve leading performance across all evaluated tasks, a result attributed primarily to the larger size of the reasoning models. STARK motivates future innovations in model architectures and reasoning paradigms for intelligent CPS by providing a structured framework to identify limitations in the spatiotemporal reasoning of LLMs and LRMs.
1 Introduction
Figure 1: Three-Tiered Architecture of STARK.
Cyber-Physical Systems (CPS) play a critical role in numerous daily and industrial applications, enabling applications like precision agriculture, environmental monitoring, and smart city infrastructure. Effective operation in these domains relies on spatiotemporal intelligence: the ability to perceive dynamic environments, understand spatiotemporal relationships, and integrate this understanding with world knowledge to make informed decisions. For instance, mobile robots need to fuse sensor data (e.g., range sensor) to navigate cluttered spaces. For various CPS applications, the common denominator is the necessity for reasoning jointly over space, time, and world knowledge.
Recent progress in large language models (LLMs) and large reasoning models (LRMs) openai2025o3o4mini ; jaech2024openai ; guo2025deepseek suggests their potential as general-purpose agents for CPS applications. Yet the community lacks rigorous benchmarks that consider the complexity in CPS applications. As shown in Figure 1, spatiotemporal reasoning in CPS typically proceeds through three stages: (1) First, state estimation. The system must perceive its environment and understand the current situation. This step involves determining what is happening, where, and when. (2) Reasoning over estimated states. Once a set of current and past states is estimated, the system needs to understand how these states interrelate across space and time. (3) Finally, world-knowledge-aware reasoning. For complex real-world scenarios, the estimated states and their spatiotemporal relationships are often interpreted with contextual information and world knowledge to make more nuanced inferences, predictions, or decisions.
However, existing benchmarks li2024stbench ; zhan2025open3dvqa ; mooney2023towards fall short of capturing this comprehensive, structured view of spatiotemporal reasoning. To address this gap, we introduce STARK (SpatioTemporal reAsoning benchmaRK), a hierarchical benchmark that measures how well models can plan and execute spatiotemporal reasoning pipelines. We designed STARK with the following features:
  • Task novelty. STARK includes under-explored problems such as field-variable prediction, spatiotemporal localization, and tracking that intertwine spatial geometry and temporal operation.
  • Task rigor. STARK requires models to craft and explain structured solution strategies. While established methods exist (e.g., multilateration), the goal of STARK is to assess whether models can identify and correctly apply the appropriate techniques for spatiotemporal tasks.
  • Multi-level evaluation. STARK is structured across three levels: (i) state estimation, (ii) reasoning over estimated states, and (iii) world-knowledge-aware reasoning that incorporates both algorithmic computation and external knowledge. This separation enables a fine-grained assessment of model capabilities across increasing levels of complexity and world-knowledge dependency.
  • Task difficulty. Unlike prior multiple-choice suites li2024stbench , STARK includes open-ended challenges: the model must output a numeric estimate given the reasoning challenges instead of picking from canned answers. Specifically, we evaluate the case where the model directly answers (DA) or invokes a Python code interpreter (CI) supplied at inference time. The latter serves as a proxy for evaluating the model’s capability of using external tools, an essential indicator of its proficiency as an AI agent. This design effectively exposes both the model’s direct reasoning ability and its competency in tool usage.
Overall, our study makes the following three primary contributions:
  • Comprehensive evaluation. We release STARK with 26 spatiotemporal reasoning scenarios and 14,552 challenge instances, providing a broad picture of LLM and LRM performance on spatiotemporal reasoning (Figure. 4). Our evaluation shows promising abilities of LRMs while also revealing their current limitations.
  • Analysis of models. We benchmark 8 LLMs and 3 LRMs, explore how direct answering (DA) and code interpreter (CI) modes affect performance, and identify limitations in spatiotemporal reasoning. Besides, we compare the models with existing first-principle baselines.
  • Open resources. All data [^1] https://huggingface.co/datasets/prquan/STARK_10k and evaluation code [^2] https://github.com/nesl/STARK_Benchmark/ are publicly available to facilitate reproducibility and foster more capable spatiotemporal reasoning systems. By exposing where LLMs and LRMs excel and fall short, STARK can inspire more effective solutions for intelligent CPS applications.
2 Related works
Benchmark Sensor type Spatial Temporal Field analysis Localization Tracking Open-end Mode
GeoQAxu2024evaluating Coordinate $\times$ $\times$ $\times$ $\times$ DA
GeoQAmooney2023towards Coordinate $\times$ $\times$ $\times$ $\times$ $\times$ DA
TemporalQAtan2023towards Text $\times$ $\times$ $\times$ $\times$ $\times$ DA
Open3DVQAzhan2025open3dvqa Image $\times$ $\times$ $\times$ $\times$ DA
STBenchli2024stbench Coordinate $\times$ $\times$ $\times$ $\times$ DA
STARK Diverse DA & CI
Table 1: Overview of existing benchmarks (DA: directly answers; CI: code interpreter)
Spatial, temporal, and spatiotemporal benchmark. GeoQA xu2024evaluating ; mooney2023towards evaluate symbolic (e.g., north-of) and geometry reasoning but ignore interaction with different sensor modalities and temporal reasoning. Temporal-QA tan2023towards ; wei2023menatqa focus on event ordering and interval logic in pure text, yet they do not include spatial constraints or open-ended numeric estimation. STBenchli2024stbench combines both spatial and temporal axes but either collapses tasks into multiple-choice or ignores other common spatial estimation tasks, such as field variable prediction, localization, or tracking. In addition, it lacks the evaluation of a spatiotemporal logic framework (Tier 2 in STARK) and reasoning with both algorithmic computation and world knowledge (Tier 3 in STARK). Vision-LLMs cheng2025v ; zhan2025open3dvqa are evaluated on spatiotemporal reasoning tasks using egocentric videos or images, focusing on object recognition, temporal event ordering, and coarse spatial grounding. However, it focuses on a single deployed sensor and lacks explicit geometric computation and multi-sensor reasoning. Lastly, none of the above works attempt to evaluate LRMs or the effectiveness of using CI.
Spatial and temporal logic frameworks. Spatial and temporal relations in STARK build directly on two classical frameworks that formally describe spatial and temporal operations. For space, we adopt the dimensionally-extended nine-intersection model (DE-9IM) clementini1993small ; clementini1994modelling implemented in ArcGIS scott2009spatial ; johnston2001using , whose topological predicates exhaustively categorize how any two geometries relate and form the backbone of modern GIS query engines, including intersects, contains, within, touches, overlaps, and their complements. For time, we rely on Allen’s interval algebra allen1983maintaining , whose 13 atomic relations provide a mutually exclusive, jointly exhaustive vocabulary for reasoning over intervals on a timeline (before, meets, overlaps, during, starts, finishes, equals and their converses). Together these frameworks constitute the symbolic operations for spatiotemporal inference. To our knowledge, no prior benchmark has directly evaluated whether LLMs can reason over both ArcGIS spatial predicates and Allen’s interval algebra in a structured setting. STARK fills this gap by incorporating both formalisms into open-ended, sensor-grounded tasks where models must apply formal spatial and temporal logic to realistic trajectories and interval structures.
3 STARK
Tier Task Sensor modality Examples
1 Spatiotemporal forecast Air quality, traffic, & temperature What is the value of sensor reading at time t? (18)
Spatiotemporal impute
Temporal impute
Spatial impute
Spatial Localization Range Where is the object most likely located? Provide the estimated [x, y] coordinates. (14)
Spatial Localization Bearing
Spatial Localization Range & bearing
Spatial Localization Region
Spatial Localization Event
Temporal Localization Event Estimate when the seismic event occurred. (15)
Spatial tracking Range Based on the available measurements at each time stamp (including the historical data), estimate the most likely [x, y] coordinates of the object in the 2D plane. (16)
Spatial tracking Bearing
Spatial tracking Range & bearing
Spatial tracking Region
Spatial tracking Event
Temporal tracking Event Based on the given TOA data, estimate when the shot was fired at each step. (17)
2 Spatial relationship Coordinate Determine whether the {geometric object 1} has the spatial relationship **{relate}** with the {geometric object 2} (19)
Temporal relationship Coordinate Determine whether the time interval {t_1} has the temporal relationship **{relate}** with the time interval {t_2} (20)
Spatiotemporal relationship Coordinate Determine whether the time interval during which the EVENT holds has the temporal relationship **{temporal_relationship}** with the reference interval {interval_2}. EVENT: {event_1} (21)
3 Landmark direction Text + Coordinate Determine whether the most accurate spatial relationship between {landmark_1} and {landmark_2} is {direction}, selecting from the options: [’north of’, ’north-east of’, ’east of’, ’south-east of’, ’south of’, ’south-west of’, ’west of’, ’north-west of’] (22)
Landmark proximity Text + Coordinate Is there a {amenity} within {distance} metres of {landmark}? (23)
Route planning Text + Coordinate Does the route from {landmark_1} to {landmark_2} pass by {landmark_3}? Route: {route} (24)
ETA calculation Text + Coordinate If I leave {landmark_1} by car at {t1}, can I arrive at {landmark_2} by {t2}? (25)
Route segment duration Text + Coordinate I drove from {landmark_1} at {t1}, and the trip took about {D} minutes. Was I likely to witness the accident that occurred along the route from {p_1} to {p_2} between {t2} and {t3}? Answer directly based on the reasoning of spatial and/or temporal information. (26)
POI prediction Text + Coordinate Determine if the user is likely to visit location {x} in the next 5 hrs? (27)
Intent prediction Text + Coordinate Determine if the user is likely to {do_X} in the next 5 hrs? (28)
Table 2: Overview of the benchmark organization
In contrast to existing benchmarks that offer fragmented views of spatiotemporal reasoning, this section introduces STARK. Table 2 summarizes tasks, sensor modalities, and example questions.
3.1 Sensor modalities
To support realistic and diverse spatiotemporal tasks in STARK, we simulate five representative sensor modalities in Figure 2: Range (distance measurement, enabling multilateration), Bearing (angular measurement, supporting triangulation), Range & Bearing (jointly output both distance and angle), Proximity (binary output for coarse localization via set intersection), and Event-based (timestamped detections for spatiotemporal inference using time-of-arrival differences). These modalities reflect common real-world sensing systems and elicit a range of geometric and combinatorial reasoning strategies. Sensor readings are injected with modality-specific noise (e.g., 1% for range sensors to mimic LiDAR accuracy; full specifications in Appendix 7.4).
3.2 Task generation
(a) Range
(b) Bearing
(c) Proximity
(d) Event (ToA)
Figure 2: State estimation. Four types of sensor modalities for localization and tracking.
(a) Reasoning over states
(b) Knowledge-aware reasoning
Figure 3: Illustration of Tier 2 and Tier 3 challenges
3.2.1 State estimation
Field variable prediction tasks aim to estimate sensor measurements across space and time, including spatial imputation (predicting a sensor’s value from neighboring sensors at the same time), temporal imputation (using historical and future values), spatiotemporal imputation (leveraging neighboring sensors’ histories), and spatiotemporal forecasting (predicting future values for neighboring sensors). We use real-world air quality data from PurpleAir noauthor_real-time_nodate and traffic data from CalTrans PeMS california_pems_nodate , and also synthesize temperature readings with controlled spatial properties.
Spatial localization aims to estimate an object’s location from sensor measurements. Each task is generated by randomly placing four sensors and a target within a $10\times 10$ grid, with all sensors sharing the same modality. Measurements are simulated (with noise) based on the target’s true position and provided to the model, which predicts the target’s $[x,y]$ location.
Temporal localization estimates the time of an event from detection times recorded by multiple spatially distributed event-based sensors. Using the same setup as spatial localization, the model’s task is to infer the event’s actual time of occurrence from these noisy arrival times.
Spatial tracking addresses dynamic object localization along continuous trajectories. In our established four-sensor, $10\times 10$ plane simulation, the object movement produces a time-stamped trajectory and sensor measurement. Inspired by SCAAT tracking welch1997scaat , we introduce partial observability via random sensor dropout. This forces the model to handle incomplete, noisy data (e.g., only two active sensors) by historical trajectory patterns and motion continuity to infer position, reflecting real-world challenges in autonomous navigation. Appendix 7.3 details trajectory simulation.
Temporal tracking estimates the timing of events occurring along a moving trajectory. We adopt the same prior sensor and environment configuration similar to SCAAT welch1997scaat .
3.2.2 Reasoning over states
Spatial relationship reasoning evaluates a model’s ability to infer qualitative spatial relationships between geometric objects. Spatial predicates defined in ArcGIS johnston2001using ; scott2009spatial lead to 35 possible combinations of geometry pairs and spatial relations (Appendix 7.5). For each instance, a pair of geometric objects is sampled and a specific relationship is generated using Shapelyshapely2007 .
Temporal interval reasoning tests a model’s ability with Allen’s interval algebra allen1983maintaining . We generate tasks by simulating two random intervals, labeling their temporal relationship, and creating a balanced negative counterpart.
Spatial-temporal reasoning evaluates a model’s ability to combine spatial and temporal logic. Using trajectories from spatial tracking, we identify intervals where a spatial predicate (e.g., Intersects, Within) holds between a trajectory and a geometry object. These event intervals are paired with reference intervals, and the model needs to decide whether a specified temporal relation (e.g., Overlaps, Meets) holds, as shown in Figure 3 (a). This task requires the model to localize events in time via spatial constraints and reason about their temporal relationships.
3.2.3 Knowledge-aware reasoning
Using ArcGIS johnston2001using ; scott2009spatial and OpenStreetMap haklay2008openstreetmap , we produce the following challenges that require world knowledge consideration: (1) Landmark direction evaluates a model’s ability to infer directional relationships between real-world landmarks. (2) Landmark proximity tests a model’s spatial reasoning in urban contexts by reasoning the presence of amenities near landmarks (e.g., Is there a police station within 500 meters of Central Park?). (3) Route planning tests whether models can connect textual travel descriptions to real-world spatial layouts. The model must decide if a given route passes by a specified landmark. (4) For ETA calculation, the model is asked to determine, given a departure time and average speed, if arrival before a specified deadline is possible based only on textual descriptions. (5) Route segment duration test the model on whether a user could plausibly be present along the segment during a given time window.
Additionally, we resort to simulating user trajectories for controlled ground-truth user intent to address data scarcity and privacy concerns in real-world mobility datasets. (6) POI prediction simulates human mobility using a finite state machine (FSM)fernandez2012human ; kerr2011activity , where each state represents a point of interest (POI). We generate a synthetic seven-day trajectory at hourly intervals. To enhance realism, we integrate contextual factors, such as weather, time, traffic, social events, etc., which are also expressed in natural language, into the models. The model must predict the next likely POI, requiring reasoning over spatiotemporal traces and language context (Figure 3 (b)). (7) Similarly, intent prediction extends POI prediction by requiring the model to infer the user’s underlying intent behind each visit. The model must predicts activity category (e.g., going home, watching a movie).
3.3 Baseline generation
We establish first-principle-based baselines for each task using theoretically grounded solutions for localization, tracking, and field variable prediction (details in Appendix 7.6). We do not construct first-principle baselines for the remaining tasks, as using spatial libraries (e.g., Shapely shapely2007 ) and geospatial services (e.g., ArcGIS johnston2001using ) yields perfect accuracy. Nonetheless, the random-guess baseline for the tasks yields 0.5 accuracy due to a balanced distribution of positive and negative pairs.
4 Experiments
Figure 4: Model ranks. Further from the center indicates better performance.
4.1 Experiment Setup
We conduct evaluations on a subset comprising 1,315 questions in STARK, due to financial constraints. For all tasks except field variable prediction, we report the Root Mean Squared Error (RMSE) against the ground truth (for both open-ended calculation and true or false questions). For field variable prediction, where outputs vary in scale, we instead report the Root Mean Squared Percentage Error (RMSPE), because of differences in data magnitude across datasets. To ensure robustness to extreme outliers, we compute all metrics using the trimmed values unless stated otherwise, i.e., excluding the top and bottom 10% of the distribution.
For evaluating the models using a code interpreter (CI), we execute the code in a sandbox environment and report the relative change in error compared to direct answering (DA). This is calculated as the percentage change in RMSE and RMSPE due to using CI (lower values indicate better performance). Code execution was supported using scipy, numpy, shapely, and filterpy libraries 2020SciPy-NMeth ; harris2020array ; shapely2007 ; filterpy . All system prompts are provided in Table 5. Costs are shown in Table 6.
4.2 Model Evaluation
model gpt-4o gpt-4.1 gpt-4.5 gpt-4o-mini Llama-4 deepseek Llama-3-8b Mistral-7B o3-mini o4-mini o3
S-impute 0.004 0.005 0.001 0.005 0.986 0.002 0.005 0.820 0.001 0.000 0.000
ST-forecast 0.098 0.075 0.098 0.109 0.100 0.083 0.143 0.070 0.152 0.106 0.092
ST-impute 0.049 0.052 0.052 0.060 0.054 0.054 0.084 0.029 0.052 0.052 0.053
T-impute 0.054 0.080 0.063 0.063 0.063 0.044 0.054 0.129 0.063 0.063 0.049
Loc-range 1.792 1.172 1.260 3.042 0.563 1.809 3.403 2.737 0.042 0.039 0.036
Loc-bearing 2.993 2.514 0.610 3.082 1.124 2.299 3.206 2.828 0.133 0.117 0.126
Loc-range-bearing 1.716 1.695 0.056 2.117 0.059 0.056 2.982 2.518 0.057 0.058 0.043
Loc-proximity 1.140 0.980 0.915 1.197 1.264 1.217 1.872 2.385 1.090 1.006 1.283
T-loc-event 0.890 0.414 0.381 0.957 0.311 0.467 0.742 2.209 0.076 0.176 0.059
S-loc-event 2.336 2.398 1.902 2.426 2.708 2.989 2.987 2.547 0.198 0.258 0.177
Track-range 2.825 3.024 2.909 3.480 2.028 2.789 4.158 3.576 0.850 0.689 0.669
Track-bearing 2.991 3.268 1.793 3.064 1.959 1.524 4.014 3.664 0.638 0.764 0.580
Track-range-bearing 2.605 2.601 0.659 2.026 1.115 0.118 4.378 3.705 0.113 0.123 0.113
Track-proximity 1.984 2.112 2.095 2.195 2.216 1.857 2.884 2.804 2.161 1.853 1.798
S-track-event 3.917 3.652 4.204 3.775 3.313 3.742 4.217 3.523 1.542 1.830 1.052
T-track-event 0.414 0.184 0.142 0.184 1.066 0.306 0.882 0.841 0.094 0.068 0.059
S relationship 0.115 0.086 0.040 0.149 0.029 0.081 0.343 0.230 0.000 0.011 0.000
T relationship 0.062 0.000 0.000 0.077 0.046 0.031 0.523 0.231 0.000 0.000 0.000
ST relationship 0.119 0.087 0.008 0.183 0.198 0.143 0.508 0.460 0.024 0.032 0.024
Landmark direction 0.200 0.200 0.200 0.333 0.000 0.333 0.667 0.600 0.133 0.000 0.000
Intent prediction 0.200 0.143 0.200 0.133 0.133 0.200 0.600 0.600 0.133 0.067 0.000
Landmark proximity 0.000 0.000 0.067 0.600 0.533 0.067 0.467 0.200 0.000 0.000 0.067
POI prediction 0.000 0.000 0.000 0.600 0.000 0.000 0.333 0.600 0.200 0.067 0.000
Route planning 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.067 0.000 0.000 0.000
ETA 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.267 0.000 0.000 0.000
Route segment 0.000 0.000 0.067 0.033 0.133 0.000 0.133 0.250 0.000 0.000 0.000
Table 3: Benchmark results. Model performance measured by RMSPE (field variable prediction) and RMSE (others). Best results are bold, second best are underlined, and third best are gray.
Table 3 and Figure 4 present the quantitative comparison between LRMs (o3, o3-mini, o4-mini) and LLMs (others). We observe that LRMs consistently outperform LLMs on Tier 1 and Tier 2 tasks, which include localization, tracking, and spatiotemporal relationship reasoning. The performance gap is especially pronounced in localization and tracking tasks, where o3 achieves 3×3$\times$3 × to 30×30$\times$30 × lower error than LLM counterparts. Among LLMs, GPT-4.5 demonstrates the strongest overall reasoning capabilities in localization and spatial/temporal relationship inference, while deepseek-chat excels in field variable prediction and tracking. Despite their strong performance in spatiotemporal reasoning, smaller LRMs such as o3-mini lag behind LLMs on Tier 3 tasks involving world knowledge, such as intent prediction and POI classification. Notably, the full o3 model maintains strong performance even in these knowledge-intensive tasks, highlighting the importance of model scale for LRMs in applying world knowledge. Since the model sizes are not disclosed, we speculate that o3 is approximately 10×10$\times$10 × larger than the o3-mini family, given the cost is also about 10×10$\times$10 × higher openai_pricing .
4.3 First-principle baselines
Figure 5: Localization (top) and tracking (bottom) compared with baselines.
Given information: Sensor location: A: [8.80, 4.06]. B: [2.97, 8.87]. C: [2.88, 8.30]. D: [2.42, 9.20]. The sensor readings are: 7.846, 6.378, 5.850, 6.554. Object location: [3.956, 5.794].
o4-mini outputted strategies 1. Write the four circle equations … 2. Subtract the “A” equation from the others to linearized … 3. Form the least-squares normal equations Write in matrix form $A\,[x\,\,y]^{\top}\approx b$ with $A=\begin{bmatrix}11.66&-9.62\\ 11.84&-8.48\\ 12.76&-10.28\end{bmatrix},\qquad b=\begin{bmatrix}-14.43\\ -10.59\\ -15.17\end{bmatrix}.$ 4. Solve the $2\times 2$ system by pseudoinverse
Multilateration baseline # Ensure ranges is a 1D array ranges = ranges.ravel() # Initial guess: centroid of the sensors initial_guess = np.mean(coordinate, axis=0) # Solve using nonlinear least squares optimization result = least_squares(residuals, initial_guess, args=(coordinate, ranges))
Figure 6: Range-based localization examples. The four sensors do not surround the target, and three are clustered together. This configuration leads to a poorly conditioned Jacobian matrix for the range equations, meaning that small errors in the sensor measurements can cause large uncertainties. Because of this, an optimization-based technique (right) may struggle to converge to a good minimum. Instead, the model chooses a Pseudo-inverse solution, aiming to mitigate the GDOP impact.
Localization.Figure 5 compares optimization-based localization approaches vs. LLMs across six tasks. Firstly, optimization-based methods generally outperform LLMs, particularly for range-only and bearing-only tasks, due to their reliance on precise geometric optimization algorithms. However, in certain instances, LRMs can match or surpass baselines. For example, in the localization range task, o3 performs the best. The baseline falters, largely due to multilateration’s sensitivity to sensor placement and non-convex optimization (Figure 6), which is exacerbated by Geometric Dilution of Precision (GDOP) when sensors are poorly arranged.
Given information: Sensor location: A: [0.35, 5.77]. B: [6.58, 8.34]. C: [5.92, 4.49]. D: [8.29, 9.01]. Step 0 | Time 0.000, sensor readings: [nan, nan, 53.6111, nan]. Step 1 | Time 0.2041, sensor readings: [nan, nan, nan, 27.3684]. Step 2 | Time 0.4082, sensor readings: [19.1300, 25.6084, nan, 25.5940]. Step 3 | Time: 0.612, sensor readings: [nan, nan, nan, 15.8711]. Step 4 | Time: 0.816, sensor readings: [nan, 19.5208, 56.7513, 12.8046]. Object location: Step 0 | [10.0, 10.0]. Step 1 | [9.796 9.796]. Step 2 | [9.592 9.592]. Step 3 | [9.388 9.388]. Step 4 | [9.184 9.184]
Outputs of o4-mini Step 0 | LRM: [6.51, 5.30] Step 1 | LRM: [5.37, 7.50] Step 2 | LRM: [4.02, 7.04] Step 3 | LRM: [2.45, 7.33] Step 4 | LRM: [8.94, 9.17]
Extended Kalman Filter (EKF) baseline Step 0 | EKF: [7.621 7.517] Step 1 | EKF: [11.040 6.512] Step 2 | EKF: [13.458 11.485] Step 3 | EKF: [16.621 12.051] Step 4 | EKF: [19.035 13.431]
Figure 7: Bearing-based tracking examples. When the initial object location is unknown and only one sensor is available at the first two steps, the LRM makes an inaccurate estimation. However, in contrast to EKF, LRM discards the inaccurate prior measurements and starts performing triangulation at the third and fifth steps.
Tracking. Figure 5 shows the performance of LLMs and Extended Kalman Filter (EKF) baselines julier2004unscented across various sensor modalities. Overall, classical EKF-based baselines outperform most LLMs by leveraging recursive filtering and explicit motion models. However, in the range and bearing tracking task, LRMs often surpass the baseline. This is mainly due to (i) the high sensitivity of bearing measurements to noise, (ii) random sensor dropout resulting in large uncertainty, and (iii) the baseline’s sensitivity to poor initialization of object location, which can cause rapid error accumulation.
As shown in Figure 7, LRMs’ consistent robustness likely stems from their ability to integrate partial observations, learn trajectory patterns, and apply contextual reasoning, maintaining accurate tracking even with ambiguous or missing data. Lastly, we observe similar results in field variable prediction tasks (Figure 9).
Given information: Sensor location: - Sensor A: [9.35, 9.99]. Sensor B: [6.75, 2.85]. Sensor C: [5.69, 5.45]. Sensor D: [5.94, 9.16] Step 0 | The corresponding sensor readings are 0.3394, 0.1314, 0.0379, 0.2101. Step 1 | The corresponding sensor readings are 0.6963, 0.7627, 0.6333, 0.5806. Step 2 | The corresponding sensor readings are 1.0990, 1.3683, 1.2629, 1.1270. Object location: Step 0 | [5.00 5.00]. Step 1 | [6.49 7.85]. Step 2 | [7.85 9.69]
Outputs of o4-mini written program Step 0 | CI: [4.82, 4.88] Step 1 | CI: [6.33, 7.83] Step 2 | CI: [-12246, 32902] Step 3 | CI: [2.45, 7.33] RMSE: 7847
Model written code def residuals(x): shooter = x[0:2] T0 = x[2] dists = np.linalg.norm(sensors_used - shooter, axis=1) t_pred = dists/v + T0 return t_pred - t_meas # Use the sensor with the earliest arrival time for an initial guess. idx_min = np.argmin(t_meas)
Figure 8: CI failure example. The code written by the model is sensitive to initialization, suggesting the need for a feedback mechanism to avoid invalid answers and improve solution quality.
4.4 Result of using a code interpreter (CI)
Table 4 summarizes the relative change in RMSE when models answer questions through Python coding with libraries compared to directly answering (DA). To mitigate extreme outliers resulting from occasional large errors produced by Python code execution, we calculated the trimmed RMSE/RMSPE for both DA and CI, removing the top and bottom 25% of data points. Several key observations emerge: (1) For localization tasks, CI notably enhances the performance of LLMs such as gpt-4o, gpt-4.1, and gpt-4o-mini, achieving error reductions of nearly 99%. (2) Conversely, for tracking tasks, CI generally degrades performance. This degradation occurs because LLM-generated tracking algorithms are vulnerable to suboptimal sensor placement, resulting in significant errors that are less frequent in DA scenarios. (3) For the LRM o3-mini, the benefits from CI are marginal overall, though occasionally coding leads to substantial increases in error (up to +354%) due to optimization-related inaccuracies. To mitigate such severe errors as shown in Figure 8, it may be necessary to implement additional mechanisms, such as guardrails with rule-based feedback and iterative refinement, to rule out the severe errors.
Table 4: Results of using CI. The table shows the relative change in RMSE compared with DA. The absolute performance is reported in Table 10. The best performance is bold, the second best is underline, and the third best is gray.
Relative change (%) gpt-4o gpt-4.1 gpt-4o-mini deepseek-chat o3-mini
S-impute -100.00 -100.00 -100.00 -100.00 -100.00
ST-forecast +91.99 +171.09 -13.94 +75.33 -7.84
ST-impute +71.60 +144.22 +136.35 +80.54 +0.00
T-impute +15.19 -16.92 +67.66 +42.51 0.00
Loc-range -98.79 -98.22 -99.08 -98.46 -12.02
Loc-bearing -84.37 -96.26 -43.51 -95.98 -11.83
Loc-range-bearing -96.75 -97.69 -97.15 -30.17 +8.70
Loc-proximity +1.00 -1.99 -6.90 -17.66 -7.49
T-loc-event -97.34 -97.52 -53.48 -95.95 -83.67
S-loc-event -91.08 -98.42 +12.03 -98.44 -80.92
Track-range +11.84 -6.23 +92.19 -0.66 +109.07
Track-bearing +21.89 -41.70 +14.25 -6.56 +68.19
Track-range-bearing -84.62 +10.74 +40.06 +2836.57 +10.02
Track-proximity +7.34 -5.91 -5.68 +27.59 -17.67
S-track-event +320.22 -59.44 +505.93 -57.34 -5.30
T-track-event -62.12 +335.23 +138.16 -60.45 +354.23
S relationship -11.50 -6.57 -14.86 -8.07 0.00
T relationship -1.54 +4.62 +1.54 +1.54 0.00
ST relationship +8.73 +8.73 +4.76 +2.65 +8.73
Landmark direction +6.67 +13.33 -$3$20.00 -13.33
Intent pred. +20.00 +25.71 +40.00 -6.67 -6.67
Landmark proximity +53.33 +40.00 +33.33 +80.00 +6.67
POI pred. 0.00 +13.33 -40.00 0.00 -20.00
Route planning 0.00 0.00 0.00 0.00 0.00
ETA calc. 0.00 0.00 0.00 0.00 0.00
Route segment dur. +60.00 +13.33 +3.33 0.00 0.00
5 Conclusion and Future Work
We present a comprehensive benchmark of 26 spatiotemporal reasoning tasks to evaluate 8 LLMs and 3 LRMs. Our findings indicate that (1) LLMs could struggle with complex geometric and sensor-driven tasks, whereas LRMs demonstrate robust performance, frequently outperforming classical first-principle baselines in localization and tracking. (2) We further explored the impact of model size on reasoning capability and observed a clear correlation. Larger LRM o3 exhibited superior performance across all tasks. Conversely, smaller LRMs such as o3-mini and o4-mini displayed limitations in integrating world knowledge, highlighting the role of model capacity in complex reasoning tasks. (3) We evaluated the impact of Code Interpreter (CI) on model performance. Interestingly, coding capabilities could enhance and diminish LLM effectiveness, suggesting the need for well-defined guardrails or rule-based constraints to guide model outputs.
Several promising avenues remain for future research. Due to benchmarking costs (Table 6), we evaluated only a subset of the full benchmark; future studies with greater resources should pursue a more exhaustive evaluation. Additionally, considering the demonstrated success of LRMs, applying reinforcement learning guo2025deepseek ; shao2024deepseekmath to fine-tune LLMs specifically for spatiotemporal reasoning tasks represents a compelling research direction. Such fine-tuning could substantially enhance the reasoning capabilities of LLMs, potentially narrowing the performance gap between general-purpose LLMs and specialized LRMs. These future directions could motivate advancements in model architectures and reasoning paradigms for intelligent CPS.
6 Acknowledgment
This research was sponsored in part by AFOSR award #FA95502210193, DEVCOM ARL award #W911NF1720196, NSF award #CNS-2325956, NIH award #P41EB028242, and Sandia National Laboratories award #2169310.